# Dipper: Diversity in Prompts for Producing Large Language Model Ensembles in Reasoning tasks

### v1

@@ Appendix C Additional Results
C.1 Performance-adjusted embedding
To study the effect of accuracy on the performance-adjusted prompt embedding matrix, we report the Spearman correlation between logdet $V$ and the ensemble performance $F(\mathcal{E})$ under different choices of $\alpha$. We observe that when $\alpha=0$, the correlation is 0.18, and it increases as $\alpha$ becomes larger. The positive correlation justifies our approach to maximize prompt diversity. The increasing correlation justifies our approach of incorporating validation accuracy into the prompt semantic embedding.
Figure 6: Spearman correlation between diversity $V$ and ensemble accuracy $F(\mathcal{E})$ on MATH.
C.2 Results on More Datasets
We also evaluate the performance of Dipper on GSM8K and MMLU-STEM. The results in Fig. 8 and Fig. 8 demonstrate that our proposed method Dipper can consistently outperform the self-ensemble baseline, further demonstrating the benefits of our method. Our proposed Dipper implementation with semantic volume optimization also consistently produces better or comparable results compared to our other variants (Random+ or Top-n) which have more unstable results, showing the usefulness of prompt diversity optimization in improving inference performance.
@@ C.2 Results on More Datasets
Figure 7: Comparison of different ensemble methods on GSM8K.
@@
Figure 8: Comparison of different ensemble methods on MMLU-STEM.

### v2

@@ Appendix C Additional Results
C.1 Results on General-Purpose Model
To show our method Dipper also generalizes to a general-purpose model (e.g., LLaMA), we evaluate its performance using LLaMA3.2-3B-it model on MATH and MMLU-STEM. The results are presented in Fig.˜10 and Fig.˜10, which demonstrate that our full Dipper implementation consistently outperforms the self-ensemble baseline and other Dipper variants across datasets.
@@ C.1 Results on General-Purpose Model
Figure 9: Comparison of different ensemble methods on MATH for the LLaMA3.2-3B-it model.
@@
Figure 10: Comparison of different ensemble methods on MMLU-STEM for the LLaMA3.2-3B-it model.
To demonstrate the generalization of our method Dipper to more recent models, we compare its variants against the baseline method on the new Qwen3-0.6B and Qwen3-1.7B models. The results are presented in Tab. 3 and 4, which suggest that our method Dipper still has the same performance advantage on recent LLMs.
Table 3: Comparison of different ensemble methods on MATH for Qwen3-0.6B
Method n=3 n=5 n=7 n=9
Self-ensemble 42.18 44.33 45.12 45.43
Dipper (Rand+) 42.18 45.03 46.31 46.98
Dipper (Top-n) 42.98 44.65 45.91 47.80
Dipper (FASV) 42.98 44.86 46.54 49.26
Table 4: Comparison of different ensemble methods on MATH for Qwen3-1.7B
Method n=3 n=5 n=7 n=9
Self-ensemble 47.38 49.54 51.07 51.70
Dipper (Rand+) 50.29 51.93 52.94 53.50
Dipper (Top-n) 51.36 53.25 53.04 53.04
Dipper (FASV) 51.57 52.62 53.25 54.51
C.2 Results on more datasets for the Qwen2-MATH-1.5B model
Apart from the MATH dataset, we also evaluate the performance of Dipper using the Qwen2-MATH-1.5B model on MMLU-STEM and GSM8K. The results in Fig.˜12 and Fig.˜12 again demonstrate that our full Dipper implementation can consistently outperform the self-ensemble baseline and achieve superior or comparable results against the other Dipper variants. The performance gains in GSM8K is more limited compared to the gains in experiments for other datasets as it is an easier dataset where the base model can already achieve high accuracy. As can be seen across all our experimental results, our full Dipper implementation comprising the theoretically-inspired semantic volume diversity optimization component achieves the most consistent performance, unlike some of the other variants.
Figure 11: Comparison of different ensemble methods on MMLU-STEM. Dipper variants outperform the self-ensemble baseline, consistent with the other experiments.
Figure 12: Comparison of different ensemble methods on GSM8K. Dipper still outperforms the self-ensemble baseline, although gains are not as obvious as in other benchmarks as it is an easier task where the base model can already perform well.
C.3 Results beyond Reasoning Tasks
To investigate the effectiveness of Dipper extending to non-reasoning tasks, we have conducted additional experiments on three challenging BIG-Bench tasks following the Instruction Induction setting from Zhou et al. (2022). As per our framework, our prompt generator first generates instruction candidates based on 5 randomly sampled demonstrations in the form of input-output pairs, before our prompt selector optimizes for the ensemble prompt composition. As shown in Tab. 5, our FASV variant consistently outperforms others, especially over the self-ensemble baseline and when the ensemble size becomes larger. This indicates that Dipper has a potential to be deployed as a general inference framework for improved performance.
Table 5: Performance Comparison on BigBench Tasks using Llama3.1-8B
Method synonyms word_unscrambling word_sorting n=3 n=5 n=7 n=9 n=3 n=5 n=7 n=9 n=3 n=5 n=7 n=9 Self-ensemble 4.38 5.0 5.88 4.88 10.25 13.0 15.25 17.38 42.125 44.25 45.25 46.0 Dipper (Rand+) 11.5 13.5 14.25 14.75 26.27 27.6 28.0 28.0 33.875 39.5 39.25 40.5 Dipper (Top-n) 10.0 13.75 12.5 17.5 22.67 24.0 25.33 25.33 28.75 28.75 36.25 33.75 Dipper (FASV) 6.25 17.5 18.75 18.75 25.33 28.0 29.33 33.33 38.75 43.75 47.5 50.0
C.4 More Results on Prompt Diversity
We also show that a strong Spearman correlation between $V$ and $F(\mathcal{E})$ exists for different datasets (e.g., GSM8K). The results presented in Fig.˜14 and Fig.˜14 demonstrate a consistent Spearman correlation between the semantic diversity $V$ and accuracy $F(\mathcal{E})$ exists. Besides, choosing different fidelity-diversity hyperparameters $\alpha$ may give different results when optimizing for the diversity.
Figure 13: Line plot showing the Spearman correlation between $V$ and $F(\mathcal{E})$ on GSM8K with different $\alpha$ values.
Figure 14: Scatter plot showing the Spearman correlation between $V$ and $F(\mathcal{E})$ on GSM8K with 2 different $\alpha$ values.
C.5 Illustration of Semantic Volume
We illustrate our semantic column metric $V$ here by comparing the semantic volume of (1) a set of 5 prompts generated by just paraphrasing a single original prompt, and (2) a set of 5 prompts randomly selected from the diverse candidate pool $\mathcal{W}$. Table˜6 shows how sets of prompts that are qualitatively observed to be more diverse have larger quantitative semantic volume.
Set of Prompts $V$
**Use a Scenario Analysis Approach**: Analyze different scenarios to determine their feasibility and impact.
*Consider Cause and Effect**: Identify potential causes and their effects to understand the question better.
*Use a Benchmarking Approach**: Compare the question to best practices or standards to find the best answer.
*Break Down the Problem**: Divide the question into smaller, manageable parts and tackle each part individually before synthesizing the overall answer.
*Apply Mathematical Logic**: Use mathematical principles and logic to solve the problem, even if it’s not a math question. -1.24
Let’s analyze this one step at a time.
Let’s break this down step by step.
Let’s tackle each part individually.
Let’s approach this incrementally.
Let’s examine this in a methodical manner. -1.80
Table 6: Example of two sets of prompts with the corresponding diversity score $V$.
C.6 Generated prompts based on 7 prompts
Below in Table˜7 we provide some examples of the generated prompts from GPT-4o based on the 7 prompts.
Prompt